In a triangle angle 1 = x - 40 ,  angle 2 =  x - 20 , angle 3 = x by 2 - 10 find the value of x

Grade:12

1 Answers

Malar Jeyaseelan
8 years ago
1 angle= x-40
2 angle=x-20
3 angle=-x/8
 
angles of a trianglle =180
 
therefore,
angle1+angle2+angle3=180
 
x-40+x-20-x/8=180
2x-x/8-60=180
15x/8=240
15x=1920
 
x=128